The following transactions of Sirta Designs took place during August 2018 in respect of kitchen stools:

Date Details Units Unit cost
01 Opening inventory 5 R50
07 Production 10 R60
19 Production 25 R72
23 Production 13 R80
29 Production 7 R90

By 31 August 2018, Rose Design had sold 50 stools.

required:

calculate the value of closing inventory, using the following inventory valuation techniques:

1. weighted average cost

2. first in, first out
Answer #1
a)Weighted Average cost:
Units available for sale= 5+10+25+13+7= 60 units
Units sold=50
Average cost per unit= cost of goods available for sale/total units available
                                            =(5*50+10*60+25*72+13*80+7*90)/60
                                            =4320/60= 72
Closing inventory= (60-50)*72= R720
b) FIFO
Units Unit cost Total
Production 23 august 3 80 240
Production 29 august 7 90 630
10 870
Closing inventory= R870